Β Show that the equation 𝑦 = π‘₯𝑒 π‘₯ is a solution of the differential equation: 𝑦 β€² βˆ’ 𝑦 βˆ’ 𝑒 π‘₯ = 0


1
Expert's answer
2022-01-18T12:02:54-0500

"(xe^x)'-xe^x-e^x=xe^x+e^x-xe^x-e^x=0"